# Control variates for variance-reduced ratio of means estimators

### v1

@@
( Preprint - October 15, 2025)
The control variates method is a classical variance reduction technique for Monte Carlo estimators that exploits correlated auxiliary variables without introducing bias. In many applications, the quantity of interest can be expressed as a ratio of expectations. We propose a variance-reduced estimator for such ratios, which applies control variates to both the numerator and the denominator. The control variate coefficients are optimized jointly to minimize the variance of the resulting estimator. This approach theoretically guarantees variance reduction and naturally extends to approximate control variates. Simulation studies show significant variance reduction, particularly when correlations between variables and control variates are strong. The practical value of the method is illustrated with a multi-fidelity aircraft design use case.
@@ 1 Introduction
An overview of common variance reduction techniques can be found in [Asmussen and Glynn, 2007]. In the setting considered here, where only a limited number of samples is available and further sampling is not possible, most methods are not directly applicable. The control variates method [Asmussen and Glynn, 2007, Sec. 2] is well-suited, as it exploits i.i.d. samples from an auxiliary variable correlated with the target variable, to reduce the variance of the Monte Carlo estimator without introducing bias. It can be seen as a form of transfer learning [Zhu et al., 2025], where information from a source variable is used to improve estimation on the target variable.
@@
One of the first references to the control variates method appears in [Hammersley and Handscomb, 1965], Section 2.6, where it is introduced under the name of “regression estimator.” The method can be interpreted as a linear regression, where the coefficient of the control variate is chosen as the slope that best explains the variability of the output through its correlation with the auxiliary variable. Later works developed this regression view further: [Lewis et al., 1987] with linear and nonlinear adjustments, [Davidson and MacKinnon, 1992] with applications to quantile estimation, and [Salaün et al., 2022] with regression-based Monte Carlo integration. Control variates have been adapted in many directions, from theoretical analyses to practical implementations; in the following, we mention a selection of approaches. [Hesterberg and Nelson, 1998] discusses the choice of control variates when estimating probabilities or quantiles. [Glynn and Szechtman, 2002] provides a comprehensive Hilbert-space perspective on control variates, establishes asymptotic optimality results, and explores their connections with other variance reduction techniques. [Hickernell et al., 2005] extends the control variates method to quasi-Monte Carlo estimators, highlighting that the optimal control variate differs from the classical Monte Carlo setting. [Müller et al., 2020] introduces Neural Control Variates, using a set of neural networks to apply control variates to parametric Monte Carlo integration. [Fox and Ökten, 2021] presents a polynomial chaos expansion method for control variates. [Maskell et al., 2022] extends the method to the estimation of constrained variables.
@@
The approximate control variates method extends the standard control variates approach to cases where the mean of the auxiliary variable is unknown, estimating it from additional samples. It is also referred to as Multi-Fidelity Monte Carlo (MFMC) [Cannamela et al., 2008] [Gorodetsky et al., 2020] [Kim et al., 2024]: the output of a computationally cheaper lower-fidelity model is used as a control variate to improve the estimation on the output of a more accurate high-fidelity model. Since both models approximate the same underlying quantity, their outputs are often highly correlated, which enables efficient variance reduction through control variates.
@@
In this work, we focus on ratio of means estimation, a quantity of interest in many applications. The standard approach consists of estimating the numerator and denominator separately by Monte Carlo, and then taking their ratio. This estimator has been analysed in [Cochran, 1977]. The control variates approach can be applied to both components to reduce variance. To the best of our knowledge, [Gordon et al., 1982] is the only work specifically extending the method to ratio of means estimation, by setting the numerator coefficient to its classical value, the one minimizing variance for a single mean, and optimizing the denominator coefficient to minimize the ratio variance.
@@
The classical coefficients for mean estimation, or those of [Gordon et al., 1982], can in fact increase the variance of the ratio of means estimator with certain correlation structures, as they are not both derived to minimize the variance of the ratio. This paper introduces new optimal coefficients that theoretically guarantee variance reduction as they are jointly optimized to minimize the variance of the estimator.
@@
The paper is organized as follows. In Section 2, the control variates method is presented for mean estimation, both in its exact and approximate forms, and for ratio of means estimation. Section 3 introduces the new optimal coefficients for variance-reduced ratio of means estimators using control variates. In Section 4, simulation results are presented, illustrating extreme correlation structures and their impact on variance reduction. Finally, in Section 5, the proposed estimators are applied to a multi-fidelity aircraft design dataset.
@@ 2.1 Control variates applied to a mean estimator
Definition 2 (Control variates estimator [Asmussen and Glynn, 2007]).
Let $(A_{i},B_{i})_{i=1\dots n}$ be i.i.d. real samples from the joint distribution of the random variables $A$ and $B$. Assume $\mathbb{E}[B]$ is known. The control variates estimator of $\mathbb{E}[A]$ is defined as
@@
The coefficient $\alpha_{c}$ needs to be estimated, which can introduce a small bias if the estimation is made from the same dataset used to compute $\widehat{A}_{CV}$. This plug-in bias [Kennedy, 2024] is generally negligible in practice, as discussed in [Owen, 2013].
@@
Definition 3 (Approximate control variates [Gorodetsky et al., 2020]).
Let $(A_{i},B_{i})_{i=1\dots n}$ be i.i.d. real samples from the joint distribution of the random variables $A$ and $B$, and let $(B_{i})_{i=n+1\dots n+m}$ be additional i.i.d. samples of $B$. The approximate control variates estimator of $\mathbb{E}[A]$ is defined as
@@
The control variate $B$ is typically chosen to mirror the variable it controls [Hesterberg and Nelson, 1998].
While the method naturally extends to multiple control variates using linear regression to estimate the various coefficients [Leluc et al., 2021], only a single control variate is considered here for clarity.
@@ 2.2 Control variates applied to ratio of means estimators
Let $(A_{i},B_{i},C_{i},D_{i})_{i=1\dots n}$ be i.i.d. real samples from the joint distribution of the random variables $A$, $B$, $C$, and $D$, where $B$ and $D$ are control variates. The CV/CV estimator of $R$ is defined as
@@
  • the coefficients introduced in [Gordon et al., 1982]:
@@
However, these choices are not always optimal: variance reduction is not guaranteed in all cases, and the variance can even be increased as illustrated analytically in A.3 and with simulations in Section 4. The variance difference between the CV/CV and MC/MC estimators can be explicitly expressed, but analysing the conditions on the covariances and variances of $A$, $B$, $C$, and $D$ that ensure variance reduction remains a delicate task.
@@ 3.1 Control variates applied to both numerator and denominator
$\text{Var}\left(\widehat{R}_{\frac{CV}{CV}}\right)-\text{Var}\left(\widehat{R}_{\frac{MC}{MC}}\right)$ $\approx-\frac{1}{n\mathbb{E}[C]^{2}}\frac{Var\bigg(\big(R\text{Cov}(B,C)-\text{Cov}(A,B)\big)D-\big(R\text{Cov}(C,D)-\text{Cov}(A,D)\big)B\bigg)}{\text{Var}(B)\text{Var}(D)-\text{Cov}(B,D)^{2}}\leq 0$ (17)
@@ 3.2 Control variates applied to the numerator only
Let $(A_{i},B_{i},C_{i})_{i=1\dots n}$ be i.i.d. real samples from the joint distribution of the random variables $A$, $B$, and $C$, and let $(B_{i})_{i=n+1\dots n+m}$ be additional i.i.d. samples of $B$. The CV/MC estimator of $R$ is defined as:
@@
The MC/MC estimator $\widehat{R}_{\frac{MC}{MC}}$ is asymptotically unbiased as $n\to\infty$ [Cochran, 1977], a property that extends to the control variates estimators $\widehat{R}_{\frac{CV}{MC}}$ and $\widehat{R}_{\frac{CV}{CV}}$ introduced previously. For smaller sample sizes, bias may arise due to the nonlinearity of the ratio form: the expectation of a ratio is generally not equal to the ratio of expectations. This bias is particularly relevant when the denominator has high variance or is near zero. Nonetheless, as illustrated in Sections 4 and 5, the impact of this bias is negligible.
@@ 4 Simulations on an academical use case
  • $(\alpha_{g},\beta_{g})$ the coefficients proposed by [Gordon et al., 1982], given in (13) and (14), where $\alpha_{g}$ minimizes the numerator variance and $\beta_{g}$ minimizes the variance of the ratio;
@@
The random variables $A,B,C,$ and $D$ are modelled as components of a four-dimensional Gaussian vector:
$(A,B,C,D)^{\top}\sim\mathcal{N}\left(\mu,\Sigma\right)\quad\text{with }\mu=\begin{pmatrix}50\\ 20\\ 10\\ 100\end{pmatrix}\quad\Sigma=\begin{pmatrix}1&\text{Cov}(A,B)&\text{Cov}(A,C)&\text{Cov}(A,D)\\ \text{Cov}(A,B)&1&\text{Cov}(B,C)&\text{Cov}(B,D)\\ \text{Cov}(A,C)&\text{Cov}(B,C)&1&\text{Cov}(C,D)\\ \text{Cov}(A,D)&\text{Cov}(B,D)&\text{Cov}(C,D)&1\end{pmatrix}$ (30)
@@
Three particular configurations are considered in the following to illustrate possible behaviours of the estimators depending on the control variates coefficients. The first corresponds to the best-case scenario for the optimal coefficients proposed in this work, i.e. the covariance structure that leads to the largest variance reduction. The second corresponds to the worst-case scenario for the coefficients of [Gordon et al., 1982], i.e. the covariance structure that leads to the largest variance increase. The last one corresponds to the best-case scenario for the coefficients of [Gordon et al., 1982]. The covariance configurations are obtained by maximizing or minimizing the variance reduction using the formulas in A.2. The optimization is carried out using the differential evolution algorithm [Storn and Price, 1997], with the objective function set to infinity whenever the tested covariance matrix is not positive definite.
@@ 4.1 Strong variance reduction
Figure 1: Boxplots of ratio of means estimations, comparing the MC/MC, CV/MC, and CV/CV estimators
Strong variance reduction with all estimators
@@ 4.2 Variance increase with non-optimal coefficients
The presented results are the worst-case scenario for the coefficients of [Gordon et al., 1982], happening with the covariance matrix
@@
Figure 2: Boxplots of ratio of means estimations, comparing the MC/MC and CV/CV estimators
Variance increase with non-optimal coefficients
While the variance reduction achieved with the optimal coefficients may not always exceed that of the other coefficients, their key advantage lies in their theoretical guarantee of variance reduction. The classical coefficients and the ones from [Gordon et al., 1982] can lead to a variance increase as illustrated in Figure 2. This result highlights the importance of jointly optimizing the coefficients to minimize the variance of the whole estimator.
4.3 Best variance reduction with the coefficients of [Gordon et al., 1982]
The presented results are the best-case scenario for the coefficients of [Gordon et al., 1982], with the following covariance matrix
@@
Figure 3: Boxplots of ratio of means estimations, comparing the MC/MC and CV/CV estimators
Best variance reduction with the coefficients from [Gordon et al., 1982] with $n=10$
@@
Figure 4: Boxplots of ratio of means estimations, comparing the MC/MC and CV/CV estimators
Equivalent variance reduction performances with $n=100$
5 Application on a multifidelity aircraft design use case
Figure 5: ONERA ALBATROS configuration with strut-braced wings (Figure 11 from [Carrier et al.,])
Figure 6: Illustrations of the $1,252$ HF and LF samples of the dataset
The CV/CV estimator is applied to estimate the strut mass fraction, defined as the ratio between the strut mass and the total mass. As illustrated in Figure 5, the strut is an external support between the fuselage and the wing that helps carry aerodynamic loads. The strut mass fraction is a quantity of interest to be optimized as it ensures the added support does not negate the aerodynamic and structural advantages of a strut-braced wing. The studied dataset contains 1,252 samples from an aircraft design study, based on [Méheut et al., 2024] which describes the multidisciplinary and multi-fidelity design process. The dataset combines outputs from high fidelity (HF) [Lannoo and Priasso, 2025] and low fidelity (LF) [Carrier et al.,] models to efficiently explore the design space. By applying the proposed CV/CV method, the strut mass fraction can be estimated more precisely with fewer HF model evaluations, which take around 2 minutes each, leveraging the LF model, which runs in only a few seconds and serves as a control variate as in MFMC.

### v2

@@
( Preprint - November 7, 2025)
The control variates method is a classical variance reduction technique for Monte Carlo estimators that exploits correlated auxiliary variables without introducing bias. In many applications, the quantity of interest can be expressed as a ratio of expectations. We propose a variance-reduced estimator for such ratios, which applies control variates to both the numerator and the denominator. The control variate coefficients are optimized jointly to minimize the variance of the resulting estimator. This approach theoretically guarantees variance reduction and naturally extends to approximate control variates. Simulation studies show significant variance reduction, particularly when correlations between variables and control variates are strong. The practical value of the method is illustrated on multi-fidelity applications: estimating a proportion in an aircraft design use case and a conditional value-at-risk in an electromagnetic dataset.
@@ 1 Introduction
An overview of common variance reduction techniques can be found in [asmussen_variance-reduction_2007]. In the setting considered here, where only a limited number of samples is available and further sampling is not possible, most methods are not directly applicable. The control variates method [asmussen_variance-reduction_2007, Sec. 2] is well-suited, as it exploits i.i.d. samples from an auxiliary variable correlated with the target variable, to reduce the variance of the Monte Carlo estimator without introducing bias. It can be seen as a form of transfer learning [zhu_recent_2025], where information from a source variable is used to improve estimation on the target variable.
@@
One of the first references to the control variates method appears in [hammersley_monte_1965], Section 2.6, where it is introduced under the name of “regression estimator.” The method can be interpreted as a linear regression, where the coefficient of the control variate is chosen as the slope that best explains the variability of the output through its correlation with the auxiliary variable. Later works developed this regression view further: [lewis_smarter_1987] with linear and nonlinear adjustments, [davidson_regression-based_1992] with applications to quantile estimation, and [salaun_regression-based_2022] with regression-based Monte Carlo integration. Control variates have been adapted in many directions, from theoretical analyses to practical implementations; in the following, we mention a selection of approaches. [hesterberg_control_1998] discusses the choice of control variates when estimating probabilities or quantiles. [glynn_new_2002] provides a comprehensive Hilbert-space perspective on control variates, establishes asymptotic optimality results, and explores their connections with other variance reduction techniques. [hickernell_control_2005] extends the control variates method to quasi-Monte Carlo estimators, highlighting that the optimal control variate differs from the classical Monte Carlo setting. [muller_neural_2020] introduces Neural Control Variates, using a set of neural networks to apply control variates to parametric Monte Carlo integration. [fox_polynomial_2021] presents a polynomial chaos expansion method for control variates. [maskell_control_2022] extends the method to the estimation of constrained variables.
@@
The approximate control variates method extends the standard control variates approach to cases where the mean of the auxiliary variable is unknown, estimating it from additional samples. It is also referred to as Multi-Fidelity Monte Carlo (MFMC) [cannamela_controlled_2008] [gorodetsky_generalized_2020] [kim_parametric_2024]: the output of a computationally cheaper lower-fidelity model is used as a control variate to improve the estimation on the output of a more accurate high-fidelity model. Since both models approximate the same underlying quantity, their outputs are often highly correlated, which enables efficient variance reduction through control variates.
@@
In this work, we focus on ratio of means estimation, a quantity of interest in many applications. The standard approach consists of estimating the numerator and denominator separately by Monte Carlo, and then taking their ratio. This estimator has been analysed in [cochran_sampling_1977]. The control variates approach can be applied to both components to reduce variance. To the best of our knowledge, [gordon_efficient_1982] is the only work specifically extending the method to ratio of means estimation, by setting the numerator coefficient to its classical value, the one minimizing variance for a single mean, and optimizing the denominator coefficient to minimize the ratio variance.
@@
The classical coefficients for mean estimation, or those of [gordon_efficient_1982], can in fact increase the variance of the ratio of means estimator with certain correlation structures, as they are not both derived to minimize the variance of the ratio. This paper introduces new optimal coefficients that theoretically guarantee variance reduction as they are jointly optimized to minimize the variance of the estimator.
@@
The paper is organized as follows. In Section 2, the control variates method is presented for mean estimation, both in its exact and approximate forms, and for ratio of means estimation. Section 3 introduces the new optimal coefficients for variance-reduced ratio of means estimators using control variates. In Section 4, simulation results are presented, illustrating extreme correlation structures and their impact on variance reduction. Finally, in Section 5.1, we demonstrate the proposed estimators on two multi-fidelity datasets: to estimate a proportion in an aircraft design dataset, and a conditional value-at-risk on an electromagnetic dataset.
@@ 2.1 Control variates applied to a mean estimator
Definition 2 (Control variates estimator [asmussen_variance-reduction_2007]).
Let $(A_{i},B_{i})_{i=1\dots n}$ be i.i.d. real samples from the joint distribution of the random variables $A,B\in L^{2}$. Assume $\mathbb{E}[B]$ is known. The control variates estimator of $\mathbb{E}[A]$ is defined as
@@
The coefficient $\alpha_{c}$ needs to be estimated, which can introduce a small bias if the estimation is made from the same dataset used to compute $\widehat{A}_{CV}$. This plug-in bias [kennedy2024semiparametric] is generally negligible in practice, as discussed in [owen_variance_2013].
@@
Definition 3 (Approximate control variates [gorodetsky_generalized_2020]).
Let $(A_{i},B_{i})_{i=1\dots n}$ be i.i.d. real samples from the joint distribution of the random variables $A,B\in L^{2}$, and let $(B_{i})_{i=n+1\dots n+m}$ be additional i.i.d. samples of $B$. The approximate control variates estimator of $\mathbb{E}[A]$ is defined as
@@
The control variate $B$ is typically chosen to mirror the variable it controls [hesterberg_control_1998].
While the method naturally extends to multiple control variates using linear regression to estimate the various coefficients [leluc_control_2021], only a single control variate is considered here for clarity.
@@ 2.2 Control variates applied to ratio of means estimators
Let $(A_{i},B_{i},C_{i},D_{i})_{i=1\dots n}$ be i.i.d. real samples from the joint distribution of the random variables $A,B,C,D\in L^{2}$, where $B$ and $D$ are control variates. The CV/CV estimator of $R$ is defined as
@@
  • the coefficients introduced in [gordon_efficient_1982]:
@@
However, these choices are not always optimal: variance reduction is not guaranteed in all cases, and the variance can even be increased as illustrated analytically in A.3 and with simulations in Section 4. The variance difference between the CV/CV and MC/MC estimators can be explicitly expressed, but analyzing the conditions on the covariances and variances of $A$, $B$, $C$, and $D$ that ensure variance reduction remains a delicate task.
@@ 3.1 Control variates applied to both numerator and denominator
$\approx-\frac{1}{n\mathbb{E}[C]^{2}}\frac{Var\bigg(\big(R\text{Cov}(B,C)-\text{Cov}(A,B)\big)D-\big(R\text{Cov}(C,D)-\text{Cov}(A,D)\big)B\bigg)}{\text{Var}(B)\text{Var}(D)-\text{Cov}(B,D)^{2}}\leq 0$ (17)
@@ 3.2 Control variates applied to the numerator only
Let $(A_{i},B_{i},C_{i})_{i=1\dots n}$ be i.i.d. real samples from the joint distribution of the random variables $A,B,C\in L^{2}$, and let $(B_{i})_{i=n+1\dots n+m}$ be additional i.i.d. samples of $B$. The CV/MC estimator of $R$ is defined as:
@@
The MC/MC estimator $\widehat{R}_{\frac{MC}{MC}}$ is asymptotically unbiased as $n\to\infty$ [cochran_sampling_1977], a property that extends to the control variates estimators $\widehat{R}_{\frac{CV}{MC}}$ and $\widehat{R}_{\frac{CV}{CV}}$ introduced previously. For smaller sample sizes, bias may arise due to the nonlinearity of the ratio form: the expectation of a ratio is generally not equal to the ratio of expectations. This bias is particularly relevant when the denominator has high variance or is near zero. Nonetheless, as illustrated in Sections 4 and 5.1, the impact of this bias is negligible.
@@ 4 Simulations on an academical use case
  • $(\alpha_{g},\beta_{g})$ the coefficients proposed by [gordon_efficient_1982], given in (13) and (14), where $\alpha_{g}$ minimizes the numerator variance and $\beta_{g}$ minimizes the variance of the ratio;
@@
The random variables $A,B,C,$ and $D$ are modeled as components of a four-dimensional Gaussian vector $(A,B,C,D)^{\top}\sim\mathcal{N}\left(\mu,\Sigma\right)$ with
$\mu=\begin{pmatrix}50\\ 20\\ 10\\ 100\end{pmatrix}\quad\Sigma=\begin{pmatrix}1&\text{Cov}(A,B)&\text{Cov}(A,C)&\text{Cov}(A,D)\\ \text{Cov}(A,B)&1&\text{Cov}(B,C)&\text{Cov}(B,D)\\ \text{Cov}(A,C)&\text{Cov}(B,C)&1&\text{Cov}(C,D)\\ \text{Cov}(A,D)&\text{Cov}(B,D)&\text{Cov}(C,D)&1\end{pmatrix}$
@@
Three particular configurations are considered in the following to illustrate possible behaviors of the estimators depending on the control variates coefficients. The first corresponds to the best-case scenario for the optimal coefficients proposed in this work, i.e. the covariance structure that leads to the largest variance reduction. The second corresponds to the worst-case scenario for the coefficients of [gordon_efficient_1982], i.e. the covariance structure that leads to the largest variance increase. The last one corresponds to the best-case scenario for the coefficients of [gordon_efficient_1982]. The covariance configurations are obtained by maximizing or minimizing the variance reduction using the formulas in A.2. The optimization is carried out using the differential evolution algorithm [storn_differential_1997], with the objective function set to infinity whenever the tested covariance matrix is not positive definite.
@@ 4.1 Strong variance reduction
Figure 1: Boxplots of ratio of means estimations, comparing the MC/MC, CV/MC, and CV/CV estimators - Strong variance reduction with all estimators
@@ 4.2 Variance increase with non-optimal coefficients
The presented results are the worst-case scenario for the coefficients of [gordon_efficient_1982], happening with the covariance matrix
@@
Figure 2: Boxplots of ratio of means estimations, comparing the MC/MC and CV/CV estimators - Variance increase with non-optimal coefficients
While the variance reduction achieved with the optimal coefficients may not always exceed that of the other coefficients, their key advantage lies in their theoretical guarantee of variance reduction. The classical coefficients and the ones from [gordon_efficient_1982] can lead to a variance increase as illustrated in Figure 2. This result highlights the importance of jointly optimizing the coefficients to minimize the variance of the whole estimator.
4.3 Best variance reduction with the coefficients of [gordon_efficient_1982]
Figure 3: Boxplots of ratio of means estimations, comparing the MC/MC and CV/CV estimator - Best variance reduction with the coefficients from [gordon_efficient_1982] with $n=10$
Figure 4: Boxplots of ratio of means estimations, comparing the MC/MC and CV/CV estimator - Equivalent variance reduction performances with $n=100$
The presented results are the best-case scenario for the coefficients of [gordon_efficient_1982], with the following covariance matrix
@@ 4.3 Best variance reduction with the coefficients of [gordon_efficient_1982]
5 Applications
5.1 Estimation of a proportion in a multifidelity aircraft design use case
The CV/CV estimator is applied to estimate the strut mass fraction, defined as the ratio between the strut mass and the total mass. The strut is an external support between the fuselage and the wing that helps carry aerodynamic loads. The strut mass fraction is a quantity of interest to be optimized as it ensures the added support does not negate the aerodynamic and structural advantages of a strut-braced wing. The studied dataset contains 1,252 samples from an aircraft design study, based on [meheut_strut_braced_2024] which describes the multidisciplinary and multi-fidelity design process. The dataset combines outputs from high fidelity (HF) [lannoo_mass_2025] and low fidelity (LF) [carrier_multidisciplinary_2022] models to efficiently explore the design space. By applying the proposed CV/CV method, the strut mass fraction can be estimated more precisely with fewer HF model evaluations, which take around 2 minutes each, leveraging the LF model, which runs in only a few seconds and serves as a control variate as in MFMC.
@@ 5.1 Estimation of a proportion in a multifidelity aircraft design use case
Figure 5: Illustrations of the $1,252$ coupled HF and LF samples of the aircraft design dataset
@@
5.2 Estimation of a conditional value-at-risk in a multi-fidelity electromagnetic dataset
A conditional Value-at-Risk (cVaR) [hong_monte_2014], also referred to as superquantile or expected shortfall [acerbi_coherence_2002] measures the average of values above a given quantile. While a quantile marks a threshold, the cVaR reflects the magnitude of extreme outcomes, providing a more robust assessment of risk widely applied in finance, reliability, and risk management [rockafellar_optimization_2000].
The cVaR $q^{*}_{\alpha}$ of $X$ of order $\alpha\in[0,1]$ is defined as
$q^{*}_{\alpha}(X)=\mathbb{E}[X|X\geq q_{\alpha}(X)]=\frac{\mathbb{E}[X\mathds{1}\{X\geq q_{\alpha}(X)\}]}{\mathbb{E}[\mathds{1}\{X\geq q_{\alpha}(X)\}]}$ (30)
where $X\in L^{2}$ is an arbitrary random variable, with cumulative distribution function $F$, and $q_{\alpha}(X)=\text{inf}\{x\in\mathbb{R}|F_{X}(x)\geq\alpha\}$ is the quantile of $X$ of order $\alpha$.
The studied multi-fidelity electromagnetic dataset is composed of $n=51$ HF samples $X^{HF}$ and $m=5,000$ additional LF samples $X^{LF}$. Aircraft certification for electromagnetic environments requires evaluating induced constraints, in terms of currents and voltages, at embedded equipment inputs. In the case of lightning indirect effects, the waveform rise time is a critical indicator, corresponding to the time in seconds between 10% and 90% of the induced current waveform.
Following the notation previously used in this paper, the cVaR estimator can be written as the ratio $R=\mathbb{E}[A]/\mathbb{E}[C]$ where
$A=X^{HF}\mathds{1}\{X^{HF}\geq q_{\alpha}(X^{HF})\}\qquad C=\mathds{1}\{X^{HF}\geq q_{\alpha}(X^{HF})\}$
The control variates are defined analogously using the LF samples:
$B=X^{LF}\mathds{1}\{X^{LF}\geq q_{\alpha}(X^{LF})\}\qquad D=\mathds{1}\{X^{LF}\geq q_{\alpha}(X^{LF})\}$
Since their exact expectations are unknown, we apply the approximate control variates method, estimating their means using the full set of $n+m=5,051$ available samples.
Due to the limited number of HF samples, bootstrapping is not appropriate, and the RVR is computed using the analytical formulas of the MC/MC and CV/CV estimators’ variance derived in Appendix A.3 in (34) and (38).
(a) Illustration of the $51$ HF and LF coupled samples of the electromagnetic dataset
(b) cVaR estimates for various quantile orders $\alpha$
(c) RVR estimates for various quantile orders $\alpha$
Figure 7: Results on the electromagnetic dataset
As shown in Figure 7(a), the HF and LF samples are strongly correlated, suggesting potential variance reduction using the CV/CV estimator. Indeed, the estimated RVR depicted in Figure 7(c) ranges from 10% to 30%. At $\alpha=0.95$, non-optimal coefficients increase variance, and the variance reduction guarantee of the optimal coefficients is illustrated.